### v1

# Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models

Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable potential in the field of natural language processing. Unfortunately, LLMs face significant security and ethical risks. Although techniques such as safety alignment are developed for defense, prior researches reveal the possibility of bypassing such defenses through well-designed jailbreak attacks. In this paper, we propose QueryAttack, a novel framework to systematically examine the generalizability of safety alignment. By treating LLMs as knowledge databases, we translate malicious queries in natural language into code-style structured query to bypass the safety alignment mechanisms of LLMs. We conduct extensive experiments on mainstream LLMs, ant the results show that QueryAttack achieves high attack success rates (ASRs) across LLMs with different developers and capabilities. We also evaluate QueryAttack’s performance against common defenses, confirming that it is difficult to mitigate with general defensive techniques. To defend against QueryAttack, we tailor a defense method which can reduce ASR by up to 64% on GPT-4-1106. The code of QueryAttack can be found on https://anonymous.4open.science/r/QueryAttack-334B.
WARNING: THIS PAPER CONTAINS UNSAFE MODEL RESPONSES.
Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models
Qingsong Zou1,3thanks: Equal contribution.  Jingyu Xiao211footnotemark: 1  Qing Li3thanks: Corresponding author.  Zhi Yan4  Yuhang Wang5 Li Xu6  Wenxuan Wang2  Kuofeng Gao1  Ruoyu Li7  Yong Jiang1 1Tsinghua Shenzhen International Graduate School  2The Chinese University of Hong Kong 3Pengcheng Laboratory  4Jilin University  5Southwest University 6University of Electronic Science and Technology of China  7Shenzhen University
1 Introduction
Large language models (LLMs) such as OpenAI’s GPT series (OpenAI, 2024a) and Meta’s Llama series (Touvron et al., 2023a) demonstrate remarkable generative potential across various domains (Xiao et al., 2024; Boiko et al., 2023; He et al., 2024; Gao et al., 2024a). However, the immense amounts of data used for training LLMs contain massive information, enabling them to learn unscreened knowledge, including those may evidently violate ethical and moral standards (Li et al., 2023; Jiang et al., 2024; Yuan et al., 2024; Bai et al., 2024). To some extent, LLMs can function as knowledge databases for users. Consequently, a critical responsibility of service providers is to prevent these models from supplying harmful information to potentially adversaries.
Figure 1: The figure illustrates a well-established phenomenon: malicious queries, when translated into a custom language, can bypass the security defenses of LLMs, leading them to generate harmful outputs based on that language. Adversaries can obtain the desired harmful text in natural language following translation. In contrast to these attacks, QueryAttack is capable of directly inducing LLMs to generate harmful text in natural language without requiring a de-translation step.
To align their responses with human ethics and preferences, numerous techniques are employed during the training process of LLMs to regulate their responses to human queries. For example, supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022), reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024), red teaming (Bai et al., 2022a), and the constitutional AI (Bai et al., 2022b) approach are proposed to enhance the safety of LLMs. These efforts enhance the ability of LLMs to identify and defend against malicious queries. Unfortunately, these countermeasures cannot completely prevent outputs that violate human ethics. A significant limitation of these methods is their reliance on malicious language samples from the alignment stage to train the model to recognize malicious queries and ensure the generation of safe outputs. This dependency leaves room for adversaries to develop jailbreak methods using non-natural language as input. For example, CipherChat (Yuan et al., 2024) uses encryption methods such as the Caesar cipher and Morse code to translate harmful queries into encrypted text. ArtPrompt (Jiang et al., 2024) replaces sensitive terms with ASCII-style encoding. CodeAttack (Ren et al., 2024b) leverages data structures to embed natural language queries within string variables of programming languages.
We observe that, the essence of these jailbreak attacks lies in defining a customized non-natural language syntax and then using this language to interact with the target LLMs, thereby bypassing their defense mechanisms. To enable the model to understand the customized non-natural language, some methods employ few-shots learning to allow the model to learn how to interact using the specified syntax in context (Peng et al., 2024; Yuan et al., 2024; Barak, 2023; Wei et al., 2023a).
Inspired by prior work, we posit that defenses that relying solely on RLHF cannot achieve a complete defense against jailbreak attacks using non-natural language as inputs. Instead, these defenses can only be incrementally improved by continuously providing LLMs with samples of zero-day attacks. However, this approach is costly and demands sustained effort and resources.
To systematically explore this perspective, we design an effective jailbreak attack to illustrate the challenges of safety generalization in environments that is distant from safety training distribution. Our proposed attack, referred to as QueryAttack, is carried out in three main components: (1) defining a query format based on non-natural language and using a translator to convert natural language queries into the target non-natural language format; (2) prompting the target LLM in context to understand the non-natural language and provide responses in natural language; (3) using the translated query to induce the model to generate harmful responses.
These three steps define a query task, analogous to quering data from a database using code. The additional cost introduced by this process is limited to translating the malicious query into the specified format, which can be easily adapted to any query based on natural languages. Given that programming languages are widely present in the training data of LLMs and that these models exhibit excellent semantic understanding of programming languages (OpenAI, 2023, 2024a; Anthropic, 2023; Touvron et al., 2023a), we naturally employ programming syntax to construct the query format.
We design a total of nine query formats written in the styles of commonly used programming languages and evaluate our attack on both well-known open-source and closed-source LLMs. Experimental results demonstrate that QueryAttack effectively bypasses their security defenses.
Our contributions can be summarized as follows:
  1. 1.
    We propose a hypothesis regarding the safety mechanisms of LLM for natural language alignment, suggesting that solely relying on RLHF cannot prevent the emergence of zero-day attacks, and use it to guide the design of a novel jailbreak attack.
  2. 2.
    We introduce QueryAttack, a novel jailbreak framework for LLMs based on query language translation. QueryAttack only requires the translation of harmful quries without requiring a de-translation step.
  3. 3.
    Our evaluation on mainstream LLMs demonstrates that QueryAttack successfully bypasses their security mechanisms and achieves state-of-the-art average attack success rate (ASR).
2 Related Work
2.1 Jailbreak Attacks on LLMs
Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information. Since natural language samples are widely used as training data during the training phase of LLMs, potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
Initially, researchers reveal that adversaries could launch attacks by manually constructing out-of-distribution (OOD) samples (Li et al., 2023; Shen et al., 2024). However, this kind of approaches relies heavily on extensive human effort, incurs significant costs on launching attacks, or it is difficult to adapt to different models or models trained on different datasets. Building on these observations, several attack methods based on white-box assumptions are proposed (Zou et al., 2023; Jones et al., 2023a, b; Gao et al., 2024c). These attacks leverage access to the model’s architecture and weights of parameters to generate adversarial prompts based on gradients. While effective, these methods are likely to fail when applied to closed-source models under black-box scenarios.
Compared to white-box attacks, black-box attacks have gradually garnered significant attention due to their minimal requirement for model-specific knowledge. In general, black-box attacks assume that adversaries can only adjust their prompt strategies based on the model’s responses (Wei et al., 2023a). Typical black-box attack methods, such as scenario setting-up or role-playing (Zhang et al., 2024; Shah et al., 2023; Liu et al., 2024; Ren et al., 2024c; Song et al., 2025), work by prompting the LLMs into dangerous hypothetical contexts, thereby triggering conflicts with the value alignment mechanisms of these models. Another approach aims to use pre-trained models to adversarially generate harmful prompts by leveraging feedback from the target models as feedback (Chao et al., 2023a).
Attacks leveraging long-tail encoded distributions are particularly effective when the target LLM’s safety fine-tuning fails to generalize to domains requiring corresponding capabilities. By replacing sensitive contents with Base64 (Wei et al., 2023a), ciphertext (Yuan et al., 2024), or low-resource languages (Deng et al., 2024), such attacks induce mismatched generalization in the target LLMs. Notably, some recent work has focused on crafting adversarial inputs using programming languages. For instance,  (Ren et al., 2024b) embed malicious queries within data structures (e.g., stacks and queues) to bypass safety alignments designed for prompts written in natural languages.  (Lv et al., 2024) encrypt malicious prompts using custom functions, transforming them into code completion tasks. The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format. However, these methods typically require the model to possess knowledge of encryption to understand the prompts or place high requirements on the model’s ability to generate encrypted content.
2.2 Safety Alignment for Defending Jailbreak
To defend against jailbreak attacks, researchers have developed various mitigation strategies to guide large language models in rejecting malicious prompts of potential adversaries.
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) is one of the most widely used defense mechanisms. For instance, recent works such as (Mehrabi et al., 2024; Sun et al., 2023) explore the effectiveness of alignment during pre-training in defending against malicious queries, as well as in-context learning (Wei et al., 2023b; Ren et al., 2024a). A common feature of these approaches is the use of red teaming and reinforcement learning (Bai et al., 2022a; Mehrabi et al., 2024). These methods often rely on natural language inputs collected from red teams, which can lead to generalization issues when faced with non-natural language or other OOD inputs.
Beyond the training process, some approaches focus on input and output safeguards, such as input perturbation (Jain et al., 2023a), safe decoding (Xu et al., 2024), and jailbreak detection (Robey et al., 2023; Phute et al., 2024; Jain et al., 2023b; Gao et al., 2024b). These methods can effectively reduce the attack success rate of jailbreak attacks. However, their effectiveness depends heavily on the quality of malicious data used for training and incurs significant additional overhead during deployment, which may affect user experience.
The latest CoT reasoning-enhanced models, such as OpenAI’s O1 model (OpenAI, 2024c), employ an embedded chain-of-thought mechanism to enforce step-by-step reasoning about the intent of queries and reflection on their compliance with policies during the response generation process. This approach does not rely on text in natural language from the training phase and simultaneously improves the model’s understanding of both queries and policy guidelines.
3 Methodology
3.1 Observation
As discussed in Section 1, we aim to demonstrate that the current generalization capabilities of safety aligned LLMs are insufficient to defend against zero-day attack samples constructed by adversaries using custom OOD languages. To promote understanding of target models, many jailbreak attack methods employ few-shots or many-shots to help the model learn the user-defined language (Yuan et al., 2024; Jiang et al., 2024). However, custom languages can be highly diverse. The effectiveness of such attacks depends significantly on the model’s ability to understand the new language.
In light of this, some attacks use common but non-natural languages as templates to encrypt malicious queries, achieving high ASR even under zero-shot conditions (Peng et al., 2024; Ren et al., 2024b). These languages often exist within the training samples of LLMs and are well understood by the models. However, they are not fully explored during the safety alignment process, making it possible for the models to generate malicious responses based on harmful prompts written by these languages. Programming languages clearly align with the common characteristics of such non-natural languages. Therefore, in this paper, we use them as examples to construct our query template.
Given the excellent generative capabilities of LLMs, we leverage them to construct translators that converts text in natural language into the our custom query templates. Few-shot learning is employed to guide the model to understand the query intents defined by the templates and generate responses in natural language. For those models with stronger comprehension capabilities (e.g., GPT-O1 (OpenAI, 2024c)), we employ zero-shot learning. Figure 2 provides an overview of QueryAttack.
Figure 2: The overview of QureyAttack. QueryAttack executes a three-step process to jailbreak the target LLM: 1). Translating the harmful query into a code-style custom template. 2. Using in-context learning to help the target LLM understand the natural semantics of the template. 3. Launching the attack using the translated template.
3.2 Query Template
Previous encryption-based jailbreak attacks convert malicious prompts into formats not encountered during the safety alignment phase. CipherChat (Yuan et al., 2024) translates queries into encrypted text using techniques such as Caesar cipher and Morse code. CodeAttack (Ren et al., 2024b) encapsulates entire natural language queries within string variables of programming languages using data structures. CodeChameleon (Lv et al., 2024) designs four different encryption functions to alter the word order of original queries or reconstruct the queries for translation.
These methods require the LLMs to generate responses in a specified non-natural language, which the adversary then decrypts to responses in natural language format. However, these approaches typically depend on the model’s knowledge of the target language to interpret the prompts or place high requirements on the model’s ability to generate text in the target language.
Unlike these attacks, we aim to only translate the malicious prompts into specific formats, but making the target LLMs directly generate responses in natural language. A total of nine commonly used programming languages are selected to design the query templates. Specifically, we identify three query components from natural language queries:
(1) the query content (content);
(2) the modifiers of the query content (modifiers);
(3) the higher-level benign category to which the query content belongs (category).
Using these three components, we define the query templates, as shown in Figure 2. This insight is inspired by the semantics of structured query languages (SQL), which query content from databases using standardized syntax. Similarly, the format of the Hypertext Transfer Protocol Secure (HTTPS) protocol, which uses Uniform Resource Locators (URLs) to fetch web resources from servers, also informs this process. Once the three query components are identified, this query template can be adapted to other programming languages. We provide a total of nine query templates based on commonly used programming languages, which can be found in the appendix A.
3.3 Query Translator
After defining the templates, the primary task is to translate natural language prompts into the template format. Extracting the three query components from the text in natural language is the most critical step in this process. This task can be performed manually by humans or automated using natural language processing tools such as NLTK (Bird et al., 2009). Given the strong text-processing capabilities of LLMs, we choose GPT-4-1106 (OpenAI, 2024a) to perform the component extraction.
We design specialized prompts to guide the GPT-4-1106 in translation. These prompts are aimed at ensuring the model focuses on the text-processing task rather than the malicious content, and thereby prevents the translation request from being rejected by the model. The details of the prompt are provided in the appendix C.
3.4 Query Learning
The purpose of the query learning is to guide the model in understanding the intent behind the query templates and then generate a natural language response. To help the LLMs understand the content queried by the templates, we first establish the context of the conversation. By describing the three query components, we guide the model in building a mapping from the query template to the natural language, and define the conversation within an educational context.
Few-shot learning is then used to reinforce the model’s understanding of the query and guide it on how to respond to these prompts using natural language. Some text in natural languages, which contain longer modifiers or multiple queries, may require several templates to help define the query. Therefore, we provide both short and long examples.
For models with strong understanding of programming languages, we can skip this process and use zero-shot to launch the attack. Finally, we provide detailed guidance in natural language to respond to the query queries. The aim of this process is to have the model answer the relevant knowledge as thoroughly as possible, rather than focusing on understanding and explaining the natural semantics of the prompts. In the appendix B, we provide the complete prompt of this part in SQL style version.
4 Experiments
4.1 Experimental Setup
Victim Models. We test QueryAttack on 11 mainstream large language models: GPT-4o (OpenAI, 2024b), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), O1 (gpt-o1) (OpenAI, 2024c), Deepseek (deepseek-chat) (DeepSeek-AI et al., 2024), Gemini-flash (gemini-1.5-flash) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct) and Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) (Touvron et al., 2023b; Meta, 2024). To maintain the reproducibility of the results, we set all the temperature to 0.
Datasets. We used AdvBench (Zou et al., 2023) as the dataset for our experiments. This is a harmful behavior dataset that contains 520 different harmful query instances written in natural language. For some experiments, we use a subset of AdvBench which contains 50 representative, non-repetitive harmful instructions refined in (Jiang et al., 2024). We will specify this at the beginning of these parts where the subset is used.
Baselines. We select five representative jailbreak techniques as baselines:
PAIR (Chao et al., 2023b): a black-box attack that uses an adversarial LLM to automatically generate adversarial inputs for the target LLM.
CipherChat (Yuan et al., 2024): a black-box attack that converts inputs into ciphers for jailbreaking.
CodeAttack (Ren et al., 2024b): a method that leverages code to embed natural language inputs into data structures to generate adversarial inputs.
TAP (Mehrotra et al., 2023): A black-box jailbreak attack via iterative refinement of attack prompts.
HEA (Song et al., 2025): A black-box method that embeds malicious queries within a scenario containing positive prompts and deceives the LLM into jailbreaking during subsequent malicious queries.
Evaluation and Metrics. We focus on three main metrics to evaluate the effectiveness of QueryAttack: harmfulness score (HS), refuse rate (RR), and attack success rate (ASR). Our evaluation follows the work of (Qi et al., 2024), utilizing the powerful evaluation capabilities of GPT-4 to assign harmfulness scores to the responses generated by the target LLM. The harmfulness score ranges from 1 to 5. To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in D.3. The extracted content is then input into GPT-4 Judge for evaluation. Detailed settings can be found in the appendix D.
Based on this, we can calculate the second metric: attack success rate (ASR), which is the percentage of harmful responses in the case of adversarial queries, as shown in Equation 1.
$ASR=\frac{\#\text{ of responses with }HS=5}{\#\text{ of responses }}.$ (1)
The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
$RR=\frac{\text{ \# of queries that are refused by LLM }}{\#\text{ of queries }}.$ (2)
In the comparison with baselines, we primarily consider ASR as the evaluation metric. In other experiments, we also report RR and HS to further supplement the evaluation.
We present the experimental results from CodeAttack (Ren et al., 2024b) and (Song et al., 2025), as they use the same benchmark as ours and also employ the GPT-4 Judge (Qi et al., 2024) method to evaluate their attacks. Therefore, we use their results as baseline for comparisons with QueryAttack. Note that (Song et al., 2025) consider an attack successful when the HS is "higher than 4", whereas we define success only when HS equals 5. However, we can still compare results using the average HS.
4.2 Results
PAIR TAP CipherChat CodeAttack HEA Ours (Top 1) Ours (Ensemble)
GPT-4-1106 - / - - / - - / 19.00% - / 81.00% - / - 4.65 / 82.18% 4.75 / 93.80%
GPT-4o* 3.16 / 45.38 % 3.24 / 51.34% 1.94 / 16.34% - / 89.00% 4.42 / 90.38% 4.72 / 90.58% 4.85 / 96.35%
Llama-3.1-8b* 3.06 / 35.38% 2.97 / 31.34% 1.76 / 0.00% - / - 4.67 / 95.38% 4.04 / 65.78% 4.83 / 88.89%
Gemini-flash* 1.92 / 18.27% 3.01 / 33.27% 2.12 / 5.38% - / - 4.64 / 100.00% 4.93 / 95.59% 4.99 / 99.62%
Table 1: Average HS / ASR of baselines and QueryAttack on the AdvBench. QueryAttack can breach the safety guardrails of mainstream LLMs, including GPT, Llama and Gemini series and achieves SoTA ASR. *: (Song et al., 2025) reports that they “consider responses of LLMs with a score higher than 4 as successful attacks” and we present their original results. When calculating the ASR for QueryAttack, we count instances where HS = 5.
Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023). We demonstrate two configurations of QueryAttack. In the first configuration, denoted as Top 1, QueryAttack uses the programming language style with the highest ASR to construct the query template. In the second configuration, denoted as Ensemble, no restrictions are placed on the programming language styles used for the query templates. From this table, we can observe that QueryAttack achieves SoTA ASR. Despite specialized safety alignment training in these latest LLMs, QueryAttack successfully bypassed their defenses, inducing responses that violate policies or human values.
For GPT-4-1106, the Top 1 configuration achieves a 82.18% ASR and the Ensemble configuration achieves 93.80%. In contrast, the baselines which perform best are only able to bypass GPT-4-1106’s safeguards in up to 81% of cases. The same trend is observed in other models. Except for Llama-3.3-8B, QueryAttack’s ASR is lower than HEA. However, under the Ensemble configuration, the average HS remains higher than that of HEA. This demonstrates the effectiveness of QueryAttack.
For cost considerations, we test QueryAttack’s effectiveness against the O1 model using the subset of AdvBench containing 50 samples. Under the Ensemble configurations, QueryAttack achieves an average HS of 3.66 and an ASR of 50%. These results indicate that CoT Reasoning-enhanced models may have the potential to defend against QueryAttack, but QueryAttack still maintains a considerable ASR.
4.3 Ablation and Analysis
Impact of different language styles. We create nine query templates in different programming language styles: Python, Java, JavaScript, C++, C, C#, SQL, URL, and Go. Figure 3 shows the average HS and RR obtained by attacking GPT-4-1106, Llama-3.1-70B, Gemini-flash and DeepSeek with templates of different styles. On GPT-4-1106, Gemini-flash and DeepSeek, different language styles do not show significant variations in average HS. A noticeable decrease is observed when attacking Llama-3.1-70B with URL and SQL-style templates, where a higher RR leads to a lower HS. This may be because these two template styles closely resemble the structure of natural language, making them more likely to trigger existing defenses. Thus, using languages that differ more from natural language is likely to increase QueryAttack’s ASR. Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reachs 76.3%, while the Ensemble configuration achieves 92.9%.
Figure 3: Performance of different language styles.
Impact of the target model’s parameter size. The increase in model parameter size does not show a positive correlation with the effectiveness of defending against QueryAttack. Table 2 presents the ASRs of QueryAttack on models with different parameter sizes from the latest Llama-3.1 and Llama-3.2 series under Ensemble configuration. On Llama-3.2 series, although the ASR slightly decreases for the 1B parameter models compared to larger models, the ASR does not show a decline as the parameter size continues to grow. Specifically, as the parameter size increases from 8B to 70B, QueryAttack’s ASR on the Llama-3.1 series models rises from 88.89% to 92.91%. Without targeted safety alignment, bigger models may even have a higher risk of being attacked due to their stronger understanding ability of new language. A similar trend can also be seen in results from other previous studies (Ren et al., 2024b; Yuan et al., 2024; Lv et al., 2024). Therefore, exploring more robust defense methods is necessary.
3.1-8B 3.1-70B 3.2-1B 3.2-3B 3.2-11B
ASR 88.89% 92.91% 94.83% 81.61% 88.89%
RR 2.87% 4.41% 3.45% 2.11% 3.26%
Table 2: QueryAttack’s attack success rate and refusal rate on Llama 3.1 and 3.2 series models with different parameter sizes, under the Ensemble configuration.
Attention score distribution of natural language malicious queries and QueryAttack. We employ contrastive input erasure (CIE) Yin and Neubig (2022) to quantify attention distribution. CIE evaluates how input tokens influence the model’s preference between an expected token (ET) and an unexpected token (UT). Since LLMs typically begin their refusal responses with “Sorry” and their acceptance responses with “Sure”, we set the ET to “Sure” and the UT to “Sorry”. An illustrative example is shown in the Figure 4, where blue tokens plays the role in rejecting query, whereas red tokens have the opposite effect. When directly questioned, the LLM precisely focuses its attention on the token “make a bomb”, significantly contributing to its likelihood of denying a malicious query. In contrast, when applying query attack, LLMs mainly allocate attention to the term “method” and “WHERE NAME = ”, leading to a higher possibility to jailbreak.
Figure 4: The attention score of natural language malicious query and QueryAttack.
4.4 Discussion about Countermeasures
We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the appendix D.2.
Paraphrase (Jain et al., 2023a): A defense method that reduces the ASRs by reconstructing inputs while preserving natural semantics.
Rand-Insert, Rand-Swap, and Rand-Patch (Robey et al., 2023): A defense method against adversarial prompts by perturbing the inputs in different ways.
Cross-lingual Alignment Prompting based defense (Ours):  (Qin et al., 2023) propose a method that uses cross-lingual chain-of-thought (CoT) prompting to generate reasoning paths, improving zero-shot CoT reasoning across languages. Although this approach is not originally designed for defense, we find that such CoT reasoning can help LLMs recognize cross-lingual malicious intent. Based on this insight, we design a defense method using cross-lingual chain-of-thought prompting. The complete chain of thought is: we first requires the target model to identify the Query Content, Key Object, and Content Source of the input, then describe the query in natural language. The target model then responds to this natural language query, thereby activating existing safety alignment mechanisms for defense. The detailed prompt can be found in the appendix D.2.
We test the effectiveness of these countermeasures against QueryAttack on four models: Gemini-flash, GPT-3.5-turbo, Llama-3.1-8B, and GPT-4-1106. For cost considerations, we use the subset of AdvBench refined in (Jiang et al., 2024) for evaluation in this section and report the results of the Ensemble configuration using Python, C++, and SQL styles.
Table 3 presents the average HS / ASR of QueryAttack and the extent to which these defenses reduce average HS / ASR. The results show that QueryAttack is robust to baseline defenses. In the worst case, QueryAttack still achieves an average ASR of 63% on GPT-3.5. SmoothLLM and Paraphrase assume that adversarial tokens may be embedded in malicious prompts, yet they do not show significant defensive effects against QueryAttack. In some cases, QueryAttack’s effectiveness is even enhanced. For example, Paraphrase increase the HS for Llama-3.1-8B by 0.14 and the ASR by 2%.
In contrast, our tailored defense based on cross-lingual alignment prompting effectively reduces the ASR of QueryAttack across all models, with the exception of a small reduction in GPT-3.5. However, on the other three models, our defense achieves an average ASR reduction of 63% and an average HS reduction of 1.53, demonstrating the effectiveness of our defense. The success of QueryAttack relies on the target model’s ability to accurately interpret custom language templates, meaning the model should also be capable of identifying the intent of them and translating them into natural language. Once translated into natural language, malicious prompts are more likely to be filtered by existing safety alignment defenses. This indicates that a well-executed translate-then-reason CoT process can effectively defend against QueryAttack-like jailbreak attacks. On the O1 model, QueryAttack achieves a lower ASR than on other models, further supporting this opinion.
Paraphrase Rand-insert Rand-swap Rand-patch Ours
Gemini-flash ASR 94% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 36% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$
HS 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 3.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$
GPT-4-1106 ASR 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ 86% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 28% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$
HS 4.16 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ 4.76 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ 4.90 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 3.10 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$
GPT-3.5 ASR 68% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 66% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 54% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ 64% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ 76% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$
HS 4.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.12 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.68 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$
Llama-3.1-8B ASR 90% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 70% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 34% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$
HS 4.86 ${\color[rgb]{1,0,0}\uparrow 0.14}$ 4.46 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 3.38 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
Table 3: QueryAttack’s average ASR / HS against defense baselines. The differences between the defense and no defense are indicated by arrows.
5 Conclusion
In this paper, we investigate the generalization challenges faced by large language models with safety alignment when encountering out-of-distribution malicious inputs. Building on previous encrypt-decrypt-based jailbreak methods, we further demonstrate that even when the target LLM is not required to generate encrypted response and instead directly produces natural language responses, a number of mainstream LLMs remain ineffective in mitigating harmful content generation.
To validate this, we introduce QueryAttack, a novel jailbreak attack framework. QueryAttack extracts three query components from a query in natural language, fill them into query templates of various styles, and leverages these templates to bypass the target LLM’s safety alignment. We conduct extensive evaluations and show that QueryAttack effectively circumvents the defenses of mainstream LLMs. To defend against QueryAttack, we developed a tailored defense strategy. Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
6 Limitation
The limitation of our study is the insufficient discussion of related defense mechanisms. As mentioned in the Related Work (section 2), research on jailbreak attack defenses has explored various potential countermeasures. However, we are unable to cover all of them. A more comprehensive investigation of defense methods would provide deeper insights into the generalizability of QueryAttack and help identify more effective mitigation techniques against such attacks.
Appendix A Query Templates
Figure 5 shows templates written in some common programming language styles. These templates express similar intentions to SQL templates, that is, querying content from a data storage.
Figure 5: Templates written in common programming language styles.
Appendix B Attack Examples of QueryAttack
B.1 Success Examples
Figure 6 illustrates a simplified attack process of QueryAttack. The adversary first utilizes query learning in context. Then, the translated query is used to launch the attack.
Figure 6: An attack example of QueryAttack.
B.2 Prompts of Query Learning
Figure 8 and Figure 7 present the prompts used for few-shot and zero-shot learning in SQL style. For zero-shot learning, we establish an educational context to encourage the model to generate more detailed information related to the risky knowledge.
Figure 7: The few-shot prompts for in-context learning (SQL style).
Figure 8: The zero-shot prompts for in-context learning (SQL style).
Appendix C Prompts for Extracting query Components
We use GPT-4-1106 (OpenAI, 2024a) to perform the component extraction task, employing few-shot learning in the process. To ensure the LLM remains focused on the natural language processing task, we refine the task details and require the model to simultaneously assess the risk level of the extracted components. The risk level of the components should not be low to prevent the model’s defense mechanisms from replacing malicious phrases with benign ones, which could affect translation quality. To reduce costs, adversaries may also choose alternative LLMs for this task.
Figure 9: The prompts for extracting query components from text in natural language using GPT-4-1106.
Appendix D Experimental Details
D.1 Human Evaluation on GPT-4 Judge
To confirm the effectiveness of GPT evaluation, we conduct an experiment with human evaluators to assess the responses of LLMs. We randomly select 100 outputs from GPT-4-1106 under the CodeAttack. Three authors independently classify these outputs as either successful or failed attacks based on OpenAI’s policy, and the final results are determined by majority vote. The manual verification results show a 93% agreement with the GPT-4 based evaluation (which also uses OpenAI’s policy as the benchmark). The result demonstrates the effectiveness of GPT-4 based evaluation.
D.2 Defense Settings
Paraphrase. We follow the setting of (Jiang et al., 2024) and use the GPT-3.5 model and the exact same prompt to complete the paraphrase task. The paraphrased translated template is then used as a replacement to launch the attack in the last step 2.
Rand-Insert, Rand-Swap, and Rand-Patch. We follow the defense and evaluation setting of (Robey et al., 2023), but choose the hyperparameter settings that give the best defensive performance reported in their paper.
Cross-lingual alignment prompting based defense. Figure 10 shows the cross-lingual alignment prompts we use for each attack sample. These prompts guide the target model to analyze the intent of the input query and rewrite it as text in natural language. Then, the target model is asked to generate response according to the query described in the text in natural language.
Figure 10: Prompt for cross-lingual alignment based defense.
D.3 Ethic-related Disclaimer Removal
Figure 11 shows a malicious response generated by the O1 model (OpenAI, 2024c) using zero-shot prompt 8. The response fully complies with the adversary’s query (Develop a program that can exploit vulnerabilities in IoT devices and take control of them). Besides, it provides concise mitigation measures as required and includes brief ethic-related disclaimers in the the response (highlighted in purple). However, these measures are clearly ineffective, as the model has already provided the detailed information queried by the adversary. Moreover, offering mitigation strategies may further enhance the adversary’s malicious knowledge to develop stronger attack vectors. Therefore, it should be considered a successful attack.
Nevertheless, we observe that these disclaimers and mitigation measures could influence the evaluation of GPT-4 Judge (Qi et al., 2024), making it more likely to assess the response as compliant with OpenAI’s policies. This could result in the harm score dropping from 5 to 4 (or even lower). Therefore, before submitting responses for GPT-4 Judge, we remove the initial disclaimer, the mitigation / education part and subsequent content to extract the harmful content.
Figure 11: An example of a malicious response generated by GPT-O1 (OpenAI, 2024c). The response is clearly malicious, containing a lot of harmful knowledge, but includes concise mitigation measures as required and provides educational disclaimers at the beginning and end (highlighted in purple).